OpenStudy (lovelyanna):

Cards with the numbers 1 through 50 are placed in a box and thoroughly shuffled. What is the probability of randomly selecting a card with a number less than 10 or a multiple of 12?

OpenStudy (lovelyanna):

@whpalmer4

OpenStudy (lovelyanna):

@jigglypuff314

OpenStudy (kropot72):

(a) How many cards have a number less than 10? (b) The multiples of 12 are found as follows: 12 * 1 = ?, 12 * 2 = ?, 12 * 3 = ?, 12 * 4 = ? and so on. How many of these multiples are less than 50? When you have the answers to (a) and (b), add them together and divide the result by 50 to find the required probability. Do you need help to do this?

OpenStudy (lovelyanna):

(a) = 1, 2, 3, 4, 5, 6, 7, 8, 9 (b) = 12, 24, 36, 48 And yes, I do, if I add all of (b), then the answer is 120, but I'm not sure if that's supposed to be correct, @kropot72

OpenStudy (lovelyanna):

Oh, and I didn't mean that the final answer is 120, by just adding (b), the answer is 120

OpenStudy (kropot72):

You have found that there are 9 cards with a number less than 10, and there are 4 cards with a number that is a multiple of 12. Therefore there are 9 + 4 = 13 cards with a number less than 10 or a multiple of 12. The required probability is therefore 13/50.

OpenStudy (lovelyanna):

Oh.. I see. Sorry, looks like I misunderstood the question. :P Thanks!

OpenStudy (kropot72):

You're welcome :)